# Mask-informed Deep Contrastive Incomplete Multi-view Clustering

### v1

@@ 3.3 Prior Knowledge-assisted Contrastive Learning
Contrastive loss: The contrastive learning approaches push the samples far away from the negative anchors while pulling in the samples with positive anchors (Chen et al., 2020; Wang & Liu, 2021). A straightforward way of constructing the positive and negative pairs is to leverage the sample connections in the graph $\hat{\mathbf{A}}$. Assuming that the neighbor points of each sample can be regarded as their positive anchors, and the points in the neighborhood of other samples can be treated as the negative anchors, we can construct the following contrastive loss to constrain the clustering distribution $\mathbf{Y}$ as formulation of InfoNCE (Yeh et al., 2022) as  [^1] The positive part is removed from denominator to alleviate the negative-positive-coupling issue.,
@@
Similarity graph $\hat{\mathbf{A}}$ measures similarities of samples as their connection possibilities in the range of $[0,1]$ as defined in (Nie et al., 2014). Simply treating the two samples with connection possibility $\hat{a}_{ij}$ larger than 0 as the positive pair is not suitable due to there is also $1-\hat{a}_{ij}$ possibility that can be regarded as negative pair. Therefore, we can further re-weight the Eq. (13) from a binary case into continual version as,
@@
3.4 Total Loss Function
@@
4 Experiments
4.1 Experimental Settings
Datasets: In our experiments, we evaluate our proposed method on eight widely used multi-view datasets, including MSRCV  [^2] https://mldta.com/dataset/msrc-v1/, CUB  [^3] https://www.vision.caltech.edu/visipedia/CUB-200.html, OutdoorScene (Hu et al., 2020), and nuswide (Zhen et al., 2019). The details descriptions of the datasets are given in the appendix.
@@ 4.1 Experimental Settings
4.2 Experimental Results
Results on complete MvC datasets: In this part, we compared the proposed method on view-complete MvC datasets with ten different approaches, including DCCA (Wang et al., 2015), AWP (Nie et al., 2018), GMC (Wang et al., 2020), OPMC (Liu et al., 2021a), MFLVC (Xu et al., 2022b), DealMVC (Chen et al., 2023), CVCL (Chen et al., 2023), EEOMVC (Wang et al., 2023a), SCMVC (Wu et al., 2024) and SCM (Luo et al., 2024). The parameters in all compared approaches are tuned with a grid search scheme as suggested in their papers to implement their best clustering performance. In Tab. 1, we present the clustering results in terms of four metrics on four MvC datasets of all methods.
From the Tab. 1, we obtain the following observations: Our method consistently performs better than other competitors on four datasets. For example, the proposed method outperforms the second performer, e.g. SCM method, with about 1.71, 0.98, 2.62, and 2.36 percentages in terms of ACC, NMI, ARI, and Fscore on the MSRCV dataset. Such advantages can be observed consistently on other datasets, which indicates the superiority of the proposed method. There are some methods, e.g. MFLVC, DealMVC, CVCL, and SCM, built upon contrastive learning. They usually treat the sample samples in different views as the positive pairs, while formulating the different samples in diverse views as the negative pairs. This type of contrastive learning manner inevitably ignores the structural information among multi-view datasets, resulting in sub-optimal clustering performance. Differently, the proposed method constructs the positive and negative pairs based on the similarities among different samples and further introduces a weighted contrastive loss to guide the model to learn more discriminative view-common feature representation for clustering. The deep MvC methods, e.g., Ours, SCM, and SCMVC perform better than traditional MvC approaches, e.g. AWP, GMC, and OPMC, demonstrating the effectiveness of deep feature representation. The above superior performance of the proposed method strongly verifies the effectiveness of the proposed method.
Results on incomplete MvC datasets: We verify the superiority of the proposed method on incomplete MvC datasets with nine different approaches: EEAE, MFAE, DCCA (Wang et al., 2015), DSIMVC (Tang & Liu, 2022), DSIMVC++ (Yan et al., 2023), CPSPAN (Jin et al., 2023), GIGA (Yang et al., 2024), MVCAN (Xu et al., 2024b) and SCM (Luo et al., 2024). The parameters in all compared approaches are tuned with a grid search scheme as suggested in their papers to implement their best clustering performance. In our experiments, we follow the strategy used in (Chen et al., 2023) to transfer the view-complete MvC data into view-complete settings with the missing rate $\eta$ in the range of $[0.1,0.3,0.5]$. In Tab. 2, we present the clustering results in terms of four metrics on four MvC datasets of all methods with missing rates varying from 0.1 to 0.5.
From the Tab. 2, we obtain the following observations: The proposed method achieves better clustering results than other compared approaches on different datasets with diverse missing rates. For example, our method performs better than the second-best results about 8.71, 3.11, 9.99, and 1.2 percentages measured by ACC, NMI, ARI, and Fscore on the OutdoorScene dataset. Such clustering performance advantages strongly indicate the effectiveness and superiority of the proposed method, which introduces a mask-informed fusion framework with a prior knowledge-assisted contrastive learning manner. The CPSPAN method imputes the view-specific latent features based on the cross-view similarity graph, which is used to find the nearest neighbor to fill missing samples. However, the imputation procedure without the true labels inevitably causes inaccurate missing view imputation, and consequentially degrades the final clustering results. GIGA and the proposed method are both imputation-free kind of approaches, which obtain better clustering results than the CPSPAN method. The MVCAN and SCM are originally designed for view-complete MvC datasets. In our experiments, we directly fill the missing samples with zeros and input MVCAN and SCM to get the clustering results. As seen from the results, MVCAN and SCM can still obtain consideration clustering performance compared with other well-designed incomplete MvC approaches in small missing ratios. The proposed method consistently performs better than MVCAN and SCM on different datasets with diverse missing ratio settings. This indicates that effectively leveraging the observation status of different samples across multiple views is beneficial for aggregating the information from IMvC data.
4.3 Model Analysis
Figure 3: The ablation studies results of the proposed method on MSRCV dataset.
Effectiveness of mask-informed fusion strategy: The mask-informed fusion strategy consists of two parts, i.e., the mask-informed feature fusion and mask-informed graph fusion, which are utilized to aggregate the incomplete multi-view information with the help of data observation status across different views. Here, we formulate two methods called wo-MFF and wo-MGF, which remove the mask-informed feature fusion and mask-informed graph fusion from the model, respectively. Due to the limited space, we only report the comparison results on MSRCV datasets in Fig. 3, and more comparison results on other datasets can be obtained in the appendix. As shown in Fig. 3, the clustering performance of the proposed method consistently outperforms than wo-MFF and wo-MGF methods, indicating the effectiveness of the proposed mask-informed fusion strategy.
Effectiveness of prior knowledge-assisted contrastive learning: The prior knowledge-assisted contrastive learning is a key component of the proposed method, which leverages the neighbor information captured from diverse views to boost the learning process of the view-common feature representation for clustering with a weighted contrastive loss. To study its effectiveness, we first remove the prior knowledge-assisted contrastive learning from our model and call it wo-WCL. As shown in Fig. 3, the clustering performance of the wo-WCL method significantly drops compared to the proposed model. In addition, we evaluate the clustering performance of the proposed model with only the prior knowledge-assisted contrastive loss without the feature reconstruction part and term it the wo-Rec method. From the results in Fig. 3, we observe that only with the prior knowledge-assisted contrastive loss can not achieve that optimal clustering performance. Thus, the feature reconstruction and prior knowledge-assisted contrastive learning are both important for the feature representation learning in the proposed method. Finally, we compare our weighted contrastive loss with the widely used contrastive loss (Chen et al., 2020) and decoupled contrastive loss (Yeh et al., 2022), which are marked as w-CL and w-DCL methods in Fig. 3. From the results, we find that 1) our weighted contrastive loss performs better than contrastive loss and decoupled contrastive loss in most cases, which indicates the weighted contrastive loss is suitable to utilize the prior knowledge of the incomplete multi-view data. 2) The weighted contrastive loss and decoupled contrastive loss achieve better results than contrastive loss indicating that the weighted contrastive loss and decoupled contrastive loss are less sensitive to the negative-positive-coupling issues (Yeh et al., 2022).
Figure 4: (a) The parameter sensitivity of the proposed method on nuswide dataset in terms of ACC, NMI, ARI, and Fscore, respectively. (b) the convergence analysis results on nuswide dataset. (c)-(d) the visualization results of the proposed method on nuswide dataset on epochs 1 and 200.
Parameter sensitivity analysis: Our proposed BMvC method consists of a key balance parameter $\lambda$ to trade-off the feature reconstruction and view-specific contrastive regularization loss. To study the parameter sensitivity, we give the clustering performance measured by ACC, NMI, ARI, and Fscore on the nuswide dataset varying with different $\lambda$ in Fig. 4, and more results are given in the appendix. From the results, we observe that the clustering performance of the proposed method slightly fluctuates with the $\lambda$. Additionally, when the parameter $\lambda$ is in the range $[10^{-}1,10^{0},10^{1}]$, the proposed method obtains considerable clustering results. Therefore, we suggest setting the parameter $\lambda$ in range $[10^{-}1,10^{0},10^{1}]$, when it is applied for some applications.
Convergence analysis and visualization: In this part, we report the clustering performance and loss function values varying from training epochs in Fig. 4. The proposed method ensures good convergence properties. Additionally, we visualize the data distribution of the view-common feature presentation under different training epochs in Figs. 4-4. From the results, we find that the proposed method can reveal more compact and clear clustering results with the training epochs increase. More convergence and visualization results can be obtained in the appendix.
5 Conclusions
In this paper, we proposed a novel incomplete multi-view clustering method equipped with a mask-informed fusion strategy and prior knowledge-assisted contrastive learning. The mask-informed fusion strategy flexibly aggregates incomplete multi-view information by considering the data observation status across multiple views, reducing the adverse impacts of missing values. The prior knowledge-assisted contrastive learning leverages the neighbor information among different views to enhance the view-common feature representation with a weighted contrastive loss. Finally, Extensive experiments on various benchmark multi-view clustering datasets on both view complete and incomplete cases were conducted to verify the efficacy of our method.
Appendix A More Results for Model Analysis
A.1 Effectiveness of mask-informed fusion strategy and prior knowledge-assisted contrastive learning
The ablation study of the proposed method in terms of ACC, NMI, ARI, and Fscore on four benchmark multi-view datasets with missing rate $\eta=0.3$ are given in Tab. 3.
@@ A.1 Effectiveness of mask-informed fusion strategy and prior knowledge-assisted contrastive learning
Table 3: The ablation study of the proposed method in terms of ACC, NMI, ARI, and Fscore on four benchmark multi-view datasets.
@@
A.2 Parameter sensitivity analysis
The parameter sensitivity analysis of the proposed method in terms of ACC, NMI, ARI, and Fscore on four benchmark multi-view datasets are given in Fig. 5.
A.3 Convergence analysis and visualization
The convergence analysis and visualization of the proposed method on four benchmark multi-view datasets are given in Figs. 6789, and 10.
@@ A.3 Convergence analysis and visualization
Appendix B Experimental Settings
B.1 Datasets
The details of four datasets are as follows:
MSRCV  [^4] https://mldta.com/dataset/msrc-v1/: This dataset contains 210 images corresponding to seven different classes. Each image is characterized by three views: 256-dimensional LBP, 512-dimensional GIST, and 210-dimensional SIFT features.
CUB  [^5] https://www.vision.caltech.edu/visipedia/CUB-200.html: This dataset consists of 600 images of different bird species, each accompanied by text descriptions, spanning 10 categories. Each sample is represented by 4096-dimensional deep image features and 300-dimensional text features.
OutdoorScene (Hu et al., 2020): This dataset includes 2688 images of outdoor scenes captured from eight different scene groups. Each image is described using four views: 512-dimensional GIST features, 432-dimensional HOG features, 256-dimensional LBP features, and 48-dimensional Gabor features.
nuswide (Zhen et al., 2019): This dataset includes 9000 images along with corresponding tags from 10 categories. Each sample is represented by 4096-dimensional deep image features and 300-dimensional text features. We randomly selected 3000 samples in our experiments.
B.2 Implementation details
The detailed structures of the view-specific encoders and decoders are respectively $D$-196-128-64 and 64-128-196-$D$, where $D$ denotes the input feature dimension. The PyTorch (Paszke et al., 2019) tool is utilized to implement the proposed method, in which the parameters are optimized with Adam optimizer (Kingma, 2014) with 0.001 initial learning rate. The proposed model is trained for 1500 epochs on all datasets, and the aggregated view-common feature representation is used to capture the cluster structures with the K-means. The features in all datasets are scaled into the range $[0,1]$ in our experiments. the trade-off parameter $\lambda$ in the proposed method is tuned in the range of $[10^{-3},10^{-2},...,10^{2},10^{3}]$. The number of neighbors in the formulated graphs is set to 15 for all datasets. All experiments are conducted on a single NVIDIA 2080Ti GPU with a Ubuntu 20.04 platform.

### v2

@@ III-C Prior Knowledge-assisted Contrastive Learning
Contrastive loss: The contrastive learning approaches push the samples far away from the negative anchors while pulling in the samples with positive anchors [35, 44]. A straightforward way of constructing the positive and negative pairs is to leverage the sample connections in the graph $\hat{\mathbf{A}}$. Assuming that the neighbor points of each sample can be regarded as their positive anchors, and the points in the neighborhood of other samples can be treated as the negative anchors, we can construct the following contrastive loss to constrain the clustering distribution $\mathbf{Y}$ as formulation of InfoNCE [45] as  [^1] The positive part is removed from denominator to alleviate the negative-positive-coupling issue.,
@@
Similarity graph $\hat{\mathbf{A}}$ measures similarities of samples as their connection possibilities in the range of $[0,1]$ as defined in [27]. Simply treating the two samples with connection possibility $\hat{a}_{ij}$ larger than 0 as the positive pair is not suitable due to there is also $1-\hat{a}_{ij}$ possibility that can be regarded as negative pair. Therefore, we can further re-weight the Eq. (13) from a binary case into continual version as,
@@
III-D Total Loss Function
@@
IV Experiments
IV-A Experimental Settings
Datasets: In our experiments, we evaluate our proposed method on four widely used multi-view datasets, including MSRCV  [^2] https://mldta.com/dataset/msrc-v1/, CUB  [^3] https://www.vision.caltech.edu/datasets/cub_200_2011/, OutdoorScene [46], and nuswide [47]. The details descriptions of the datasets are given in follows:
MSRCV: This dataset contains 210 images corresponding to seven different classes. Each image is characterized by three views: 256-dimensional LBP, 512-dimensional GIST, and 210-dimensional SIFT features.
CUB: This dataset consists of 600 images of different bird species, each accompanied by text descriptions, spanning 10 categories. Each sample is represented by 4096-dimensional deep image features and 300-dimensional text features as doing in [47].
OutdoorScene: This dataset includes 2688 images of outdoor scenes captured from eight different scene groups. Each image is described using four views: 512-dimensional GIST features, 432-dimensional HOG features, 256-dimensional LBP features, and 48-dimensional Gabor features.
nuswide: This dataset includes 9000 images along with corresponding tags from 10 categories. Each sample is represented by 4096-dimensional deep image features and 300-dimensional text features as doing in [47]. We randomly selected 3000 samples in our experiments.
@@ IV-A Experimental Settings
TABLE III: The ablation study of the proposed method in terms of ACC, NMI, ARI, and Fscore on four benchmark multi-view datasets.
@@
IV-B Implementation details
The detailed structures of the view-specific encoders and decoders are respectively $D$-196-128-64 and 64-128-196-$D$, where $D$ denotes the input feature dimension. The PyTorch [48] tool is utilized to implement the proposed method, in which the parameters are optimized with Adam optimizer [49] with 0.001 initial learning rate. The proposed model is trained for 1500 epochs on all datasets, and the aggregated view-common feature representation is used to capture the cluster structures with the K-means. The features in all datasets are scaled into the range $[0,1]$ in our experiments. the trade-off parameter $\lambda$ in the proposed method is tuned in the range of $[10^{-3},10^{-2},...,10^{2},10^{3}]$. Temperature $\tau$ is set to 1. The number of neighbors in the formulated graphs is set to 15 for all datasets. All experiments are conducted on a single NVIDIA 2080Ti GPU with a Ubuntu 20.04 platform.
IV-C Experimental Results
Results on complete MvC datasets: In this part, we compared the proposed method on view-complete MvC datasets with ten different approaches, including DCCA [50], AWP [51], GMC [52], OPMC [53], MFLVC [38], DealMVC [54], CVCL [54], EEOMVC [55], SCMVC [56] and SCM [40]. The parameters in all compared approaches are tuned with a grid search scheme as suggested in their papers to implement their best clustering performance. In Tab. I, we present the clustering results in terms of four metrics on four MvC datasets of all methods.
From the Tab. I, we obtain the following observations: Our method consistently performs better than other competitors on four datasets. For example, the proposed method outperforms the second performer, e.g. SCM method, with about 1.71, 0.98, 2.62, and 2.36 percentages in terms of ACC, NMI, ARI, and Fscore on the MSRCV dataset. Such advantages can be observed consistently on other datasets, which indicates the superiority of the proposed method. There are some methods, e.g. MFLVC, DealMVC, CVCL, and SCM, built upon contrastive learning. They usually treat the sample samples in different views as the positive pairs, while formulating the different samples in diverse views as the negative pairs. This type of contrastive learning manner inevitably ignores the structural information among multi-view datasets, resulting in sub-optimal clustering performance. Differently, the proposed method constructs the positive and negative pairs based on the similarities among different samples and further introduces a weighted contrastive loss to guide the model to learn more discriminative view-common feature representation for clustering. The deep MvC methods, e.g., Ours, SCM, and SCMVC perform better than traditional MvC approaches, e.g. AWP, GMC, and OPMC, demonstrating the effectiveness of deep feature representation. The above superior performance of the proposed method strongly verifies the effectiveness of the proposed method.
Results on incomplete MvC datasets: We verify the superiority of the proposed method on incomplete MvC datasets with nine different approaches: EEAE, MFAE, DCCA [50], DSIMVC [57], DSIMVC++ [58], CPSPAN [59], GIGA [23], MVCAN [60] and SCM [40]. The parameters in all compared approaches are tuned with a grid search scheme as suggested in their papers to implement their best clustering performance. In our experiments, we follow the strategy used in [54] to transfer the view-complete MvC data into view-complete settings with the missing rate $\eta$ in the range of $[0.1,0.3,0.5]$. We repeat all approaches for 5 times and present the mean clustering results in terms of four metrics on four MvC datasets of all methods with missing rates varying from 0.1 to 0.5 in Tab. II.
From the Tab. II, we obtain the following observations: The proposed method achieves better clustering results than other compared approaches on different datasets with diverse missing rates. For example, our method performs better than the second-best results about 8.71, 3.11, 9.99, and 1.2 percentages measured by ACC, NMI, ARI, and Fscore on the OutdoorScene dataset. Such clustering performance advantages strongly indicate the effectiveness and superiority of the proposed method, which introduces a mask-informed fusion framework with a prior knowledge-assisted contrastive learning manner. The CPSPAN method imputes the view-specific latent features based on the cross-view similarity graph, which is used to find the nearest neighbor to fill missing samples. However, the imputation procedure without the true labels inevitably causes inaccurate missing view imputation, and consequentially degrades the final clustering results. GIGA and the proposed method are both imputation-free kind of approaches, which obtain better clustering results than the CPSPAN method. The MVCAN and SCM are originally designed for view-complete MvC datasets. In our experiments, we directly fill the missing samples with zeros and input MVCAN and SCM to get the clustering results. As seen from the results, MVCAN and SCM can still obtain consideration clustering performance compared with other well-designed incomplete MvC approaches in small missing ratios. The proposed method consistently performs better than MVCAN and SCM on different datasets with diverse missing ratio settings. This indicates that effectively leveraging the observation status of different samples across multiple views is beneficial for aggregating the information from IMvC data.
IV-D Model Analysis
Effectiveness of mask-informed fusion strategy: The mask-informed fusion strategy consists of two parts, i.e., the mask-informed feature fusion and mask-informed graph fusion, which are utilized to aggregate the incomplete multi-view information with the help of data observation status across different views. Here, we formulate two methods called wo-MFF and wo-MGF, which remove the mask-informed feature fusion and mask-informed graph fusion from the model, respectively. As shown in Tab. III, the clustering performance of the proposed method consistently outperforms than wo-MFF and wo-MGF methods, indicating the effectiveness of the proposed mask-informed fusion strategy.
Effectiveness of prior knowledge-assisted contrastive learning: The prior knowledge-assisted contrastive learning is a key component of the proposed method, which leverages the neighbor information captured from diverse views to boost the learning process of the view-common feature representation for clustering with a weighted contrastive loss. To study its effectiveness, we first remove the prior knowledge-assisted contrastive learning from our model and call it wo-WCL. As shown in Tab. III, the clustering performance of the wo-WCL method significantly drops compared to the proposed model. In addition, we evaluate the clustering performance of the proposed model with only the prior knowledge-assisted contrastive loss without the feature reconstruction part and term it the wo-Rec method. From the results in Tab. III, we observe that only with the prior knowledge-assisted contrastive loss can not achieve that optimal clustering performance. Thus, the feature reconstruction and prior knowledge-assisted contrastive learning are both important for the feature representation learning in the proposed method. Finally, we compare our weighted contrastive loss with the widely used contrastive loss [35] and decoupled contrastive loss [45], which are marked as w-CL and w-DCL methods in Tab. III. From the results, we find that 1) our weighted contrastive loss performs better than contrastive loss and decoupled contrastive loss in most cases, which indicates the weighted contrastive loss is suitable to utilize the prior knowledge of the incomplete multi-view data. 2) The weighted contrastive loss and decoupled contrastive loss achieve better results than contrastive loss indicating that the weighted contrastive loss and decoupled contrastive loss are less sensitive to the negative-positive-coupling issues [45].
@@ IV-D Model Analysis
Parameter sensitivity analysis: Our proposed BMvC method consists of a key balance parameter $\lambda$ to trade-off the feature reconstruction and view-specific contrastive regularization loss. To study the parameter sensitivity, we give the clustering performance measured by ACC, NMI, ARI, and Fscore on four datasets varying with different $\lambda$ in Fig. 3, and more results are given in the appendix. From the results, we observe that the clustering performance of the proposed method slightly fluctuates with the $\lambda$. Additionally, when the parameter $\lambda$ is in the range $[10^{-}1,10^{0},10^{1}]$, the proposed method obtains considerable clustering results. Therefore, we suggest setting the parameter $\lambda$ in range $[10^{-}1,10^{0},10^{1}]$, when it is applied for some applications.
Convergence analysis and visualization: In this part, we report the clustering performance and loss function values varying from training epochs on different datasets in Fig. 4. The proposed method ensures good convergence properties. Additionally, we visualize the data distribution of the view-common feature presentation under different training epochs in Figs. 567, and 8. From the results, we find that the proposed method can reveal more compact and clear clustering results with the training epochs increase.
V Conclusions
In this paper, we proposed a novel incomplete multi-view clustering method equipped with a mask-informed fusion strategy and prior knowledge-assisted contrastive learning. The mask-informed fusion strategy flexibly aggregates incomplete multi-view information by considering the data observation status across multiple views, reducing the adverse impacts of missing values. The prior knowledge-assisted contrastive learning leverages the neighbor information among different views to enhance the view-common feature representation with a weighted contrastive loss. Extensive experiments on various benchmark multi-view clustering datasets on both view complete and incomplete cases were conducted to verify the efficacy of our method.
The method demonstrates a slight advantage over others at missing rates of 0.5. This degradation occurs because views with high miss ratios make it challenging to construct accurate neighborhoods for the proposed prior knowledge-assisted contrastive loss, leading to performance degradation. In contrast, at missing rates between 0 and 0.3, our method shows significant advantages, indicating its suitability for incomplete multi-view datasets with lower missing rates. In future research, we propose to expand our approach to better address challenges associated with extreme missing rates.
Acknowledgment
The authors wish to gratefully acknowledge the anonymous reviewers for the constructive comments of this paper.